the area of a rectangular room is 84sq m. if the breath of the room is 5m less than its length, find the perimeter of the room
Answer by checkley79(3341)

A=84=LW OR W=84/L
P=2L+2W
W=L-5
84/L=L-5
L(L-5)=84
L^2-5L-84=0
(L-12)(L+7)=0
L-12=0
L=12 ANS. FOR THE LENGTH.
W=84/712
W=7 ANS. FOR THE WIDTH.
PROOF:
84=7*12
84=84
P=2*7+2*12
P=14+24=38 ANS. FOR THE PERIMETER.